OpenStudy (anonymous):

Which of the following ordered pairs represents the solution to the system given below? x − 4y = 7 5x + 9y = 6 (3, −1) (3, 1) (1, −3) (−1, −3)

OpenStudy (anonymous):

1,-3

OpenStudy (carolinar7):

no -3,1

OpenStudy (anonymous):

A student is trying to solve the system of two equations given below: Equation P: x + y = 8 Equation Q: 2x + 5y = 24 Which of the following steps can be used to eliminate the x term? 2(x + y = 8) −2(x + y = 8) −1(2x + 5y = 24) −2(2x + 5y = 24)

OpenStudy (carolinar7):

I mean 3,-1

OpenStudy (anonymous):

A student is trying to solve the system of two equations given below: Equation P: x + y = 8 Equation Q: 2x + 5y = 24 Which of the following steps can be used to eliminate the x term? 2(x + y = 8) −2(x + y = 8) −1(2x + 5y = 24) −2(2x + 5y = 24)